# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v3

@@ Abstract
For random-walk Metropolis (RWM) and parallel tempering (PT) algorithms, an asymptotic acceptance rate of around 0.234 is known to be optimal in certain high-dimensional limits. However, its practical relevance is uncertain due to restrictive derivation conditions. We synthesise previous theoretical advances in extending the 0.234 acceptance rate to more general settings, and demonstrate its applicability with a comprehensive empirical simulation study on examples examining how acceptance rates affect Expected Squared Jumping Distance (ESJD). Our experiments show the optimality of the 0.234 acceptance rate for RWM is surprisingly robust even in lower dimensions across various proposal, multimodal distributions that may not have an i.i.d. product density, and curved Rosenbrock target distributions with nonlinear correlation structure. Parallel tempering experiments also show that the idealized 0.234 spacing of inverse temperatures may be approximately optimal for low dimensions and non i.i.d. product target densities, and that constructing an inverse temperature ladder with spacings given by a swap acceptance of 0.234 is a viable strategy.
@@ 1 Introduction
[29] proved an important result for independently and identically distributed (i.i.d.) product target densities: in the state space high-dimensional limit $d\to\infty$, the asymptotic acceptance rate of a Metropolis algorithm is central to its efficiency. Under their assumptions, an asymptotic acceptance rate of approximately 0.234 maximises the efficiency of a random-walk Metropolis (RWM) algorithm with a Gaussian increment proposal distribution starting at stationarity. However, their assumptions on the target and proposals are very restrictive. Many practical scenarios often involve finite and relatively small dimensions. Additionally, many probability distributions do not have densities of the i.i.d. product form, and for those that do, sampling a one-dimensional target is much simpler than using RWM. A burgeoning active research area is dedicated to showing that this figure still applies beyond these assumptions, and much of this focuses on the i.i.d. product assumption.
Many papers have shown that the optimal average acceptance rate of 0.234 for RWM holds greater generality than the assumptions used in Roberts et al. [29]. For more general theoretical results, Sherlock [38] replaces the restrictive i.i.d. product target and Gaussian proposal setting of Roberts et al. [29] with a set of “shell”, relative variability, and eccentricity conditions on the target, eccentricity condition on the jump distance matrix, and shell condition on the proposal. This framework does not require a product form and is based on the geometric properties of the (log) target distribution in high dimensions. Yang et al. [42] makes the 0.234 result applicable to targets with sparse dependency structures, such as those arising from Bayesian graphical models, using a set of sufficient conditions based on local dependencies and bounded asymptotic behaviour of derivatives. We expand on these results in Section 2. The 0.234 acceptance rate in optimal scaling has also been shown to apply to the spacing and swapping between inverse temperatures in the parallel tempering MCMC method [1, 36]. That being said, there are some cases in which 0.234 is not optimal, such as in the case of discontinuous targets [25] where the optimal acceptance rate is approximately 0.1353. Additionally, these results of optimality are still theoretical ideals, with required assumptions that could be further relaxed, and in the limit of $d\to\infty$ where the Markov chain converges to a diffusion process.
@@
These results are of great interest to practitioners who would like to understand how to tune their MCMC algorithms to maximise their efficiency. Previous simulation studies have been carried out on more general target distributions like products of independent, non-identical components [7, 33], and for the Markov modulated Poisson process [37], but there is still a lack of thorough experimentation and guidance for how handy the 0.234 rule is on realistic target distributions that may have lower or higher dimensionality and/or multimodality. Our research examines how necessary these assumptions are for the theoretical result to remain relevant: where can we relax some assumptions in lower dimensions and still have the optimal acceptance rate of approximately 0.234, and where can we not do this? And in the end, how applicable is the 0.234 figure for the practitioner, who may want to use MCMC for complicated target distributions beyond the necessary assumptions in finite dimensions? We thoroughly dissect various aspects of the Metropolis algorithm and experiment with them, and in doing so, we show empirically where the theoretical ideal value can still align with more realistic scenarios.
@@ 2.1 Metropolis Algorithm
A Metropolis algorithm [21] constructs a Markov chain $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots$. If the chain is constructed to be irreducible and aperiodic, it is guaranteed to have the target distribution $\pi$ as its unique stationary distribution. This means that as the number of steps $t$ becomes large, the distribution of $\mathbf{x}^{(t)}$ converges to $\pi$. Therefore, after an initial burn-in period, the subsequent states of this single, long chain can be used as a sample from the target distribution $\pi$. When taking each step in the Markov chain, we generate a new state $\mathbf{y}=\mathbf{x}^{(t)}+\epsilon$ where $\epsilon$ is sampled from an increment “proposal” density $Q$ having $Q(\epsilon)=Q(-\epsilon)$ for all $\epsilon$ in its domain and accept this new state with probability $\alpha(\mathbf{x},\mathbf{y})$ given by $\alpha=\min\left\{1,\dfrac{\pi(\mathbf{y})}{\pi(\mathbf{x})}\right\}$, where, by a common abuse of notation, we let $\pi(\cdot)$ also denote the probability density function of the target distribution $\pi$. If the proposed value is accepted, we set the next state $\mathbf{x}^{(t+1)}=\mathbf{y}$. Otherwise, we set $\mathbf{x}^{(t+1)}=\mathbf{x}^{(t)}$.
@@ 2.2 Optimal Scaling Framework
Optimal Scaling [29, 33] is one of the most successful frameworks for performing asymptotic analysis of high-dimensional MCMC methods, and provides mathematically-grounded guidance on how to best optimise MCMC performance by tuning the “scaling” parameter(s) of the proposal distribution. Typically, a multivariate Gaussian distribution with mean 0 is used as the increment proposal distribution. The scaling of the Gaussian proposal is then given by a scaling factor $\sigma$, and this is often used as a variance $\sigma^{2}$ that is applied to the $d\times d$ identity matrix $I_{d}$. So, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{d})$. If the proposal’s variance is too small, many proposed steps are accepted by the algorithm, but each step does not explore the state space much. Vice versa, if the proposal’s variance is too large, too few proposed steps are accepted, leading to slow exploration of the state space since the algorithm stays at a state for too long. Therefore, the proposal’s variance is crucial to the performance of the algorithm; the optimal scaling should balance the exploration of new areas in the state space with exploiting high-density areas of the target distribution. Different Metropolis algorithms have shown to have different optimal choices of the proposal distribution variance, $\sigma^{2}$.
@@
where $n$ is the total number of iterations of the algorithm. Maximising the ESJD aligns with minimizing the first-order auto-correlation of the Markov chain and subsequently maximises efficiency if the higher-order auto-correlations are monotonically increasing relative to the first-order auto-correlation [26, 42]. It is worth mentioning that ESJD primarily measures local convergence performance. Because heavy-tailed targets disrupt the monotonic relationship between first-order and higher-order autocorrelations, achieving global convergence for heavy-tailed targets requires maximising an adapted version of ESJD [17]. However, we do not consider heavy-tailed targets in this work, so the working definition of ESJD is sufficient.
There are other notions of efficiency of a Markov chain, but in the high-dimensional limit $d\to\infty$, if the moment conditions in Assumption 2.1 holds, the chain converges to a diffusion process, and all efficiency measures are effectively equivalent [29, 33]. A key result of these referenced papers is that, given a Metropolis algorithm with a Gaussian increment proposal distribution $Q=\mathcal{N}(0,\frac{\ell^{2}}{d}I_{d})$ where $\ell>0$ is a fixed scaling constant and $I_{d}$ is the identity matrix, maximising the diffusion’s speed measure $h(\ell)$, which is a function of the scaling constant $\ell$, yields the most efficient asymptotic diffusion. Furthermore, the speed measure has a clear relation to a much simpler quantity to tune for: the asymptotic acceptance rate of the proposed new states (moves) of the algorithm, defined as
@@
Under the conditions in Assumption 2.1, Roberts et al. [29] showed that for an RWM algorithm with a Gaussian proposal distribution starting in stationarity, the sped-up process converges to a limiting Langevin diffusion. Maximizing the speed of this limiting process, in turn, corresponds to tuning the algorithm to an asymptotic acceptance rate of approximately 0.234.
@@
(Sherlock [38]) Let $\mathbf{x}$ be a random variable from the target density $\pi$, $M(\mathbf{x})=||\nabla\log\pi(\mathbf{x})||$ be the norm of the gradient of the log target, and $H(\mathbf{x})=-\nabla^{2}\log\pi(\mathbf{x})$ be the negative Hessian matrix of the log target. The primary conditions are:
@@
Put together, these geometric regularities recover the same weak-convergence proof as Roberts et al. [29] but with weaker assumptions on the target.
@@ 3 Random-Walk Metropolis Simulations
In this section, we explore the applicability of the 0.234 acceptance rate figure for RWM algorithms using several practical RWM examples in lower dimensions. We begin with simple examples where the target densities meet all the key conditions in Roberts et al. [29] (provided as Assumption 2.1) required for the 0.234 acceptance rate to be optimal. Next, we investigate examples that do not use Gaussian proposal distributions, and then we explore other target densities that may lack these assumptions, such as smoothness or the i.i.d. product form. We estimate the Expected Squared Jumping Distance (ESJD) of the algorithm as a metric of the algorithm’s efficiency as mentioned in Section 2.2 and aim to see under what conditions an acceptance rate of approximately 0.234 maximises the ESJD for various target and proposal distributions in lower dimensions.
@@
The first example sets $f_{1}$ with shape parameter 3 and scale parameter 2, and the second example sets $f_{2}$ with shape parameters 3 and 2. We use a multivariate Gaussian with a mean of 0 as the increment proposal distribution in both cases. We run this experiment in dimensions $d\in\{2,5,10,30,50,100\}$.
@@
Since most ESJD versus acceptance rate plots look fairly similar and have a similar progression as dimension changes, the full set of curves for ESJD versus acceptance rate plots for other subsections are in Appendix A, including the results for the i.i.d. Gamma in Figure 9. As an aside, because the plateau widens appreciably in higher dimensions, the coarse 40-point grid pinpoints the optimum to approximately ±$\pm 0.01$.
@@ 3.2 Non-Gaussian Proposal Densities
Next, we investigate the efficacy of different proposal distributions for the Metropolis algorithm. One of the conditions assumed by the 0.234 theorem of [29] is the use of a multivariate Gaussian increment proposal distribution $Q=\mathcal{N}(0,\sigma^{2}I_{d})$. The additional results in [33] depend on the assumptions of light-tailed proposals. In contrast, [16] extend this analysis by examining the behaviour of the algorithms with heavy-tailed proposals. They find that, unlike the diffusion behaviour seen with light-tailed proposals, heavy-tailed proposals lead to a different dynamic characterized by abrupt movements followed by periods of inactivity. Furthermore, [24] investigated optimal scaling results for heavy-tailed proposal distributions such as the Cauchy distribution, and spherical proposal distributions of fixed radius. In their RWM simulation study with three different continuous i.i.d. product target densities, both the Gaussian and Cauchy proposals had an estimated asymptotically optimal acceptance rate of approximately 0.234 in $d=100$, and a spherical proposal with a fixed radius of 2.38 achieved an optimal acceptance rate of about 0.234 in $d=2$ and $d\geq 10$. Finally, Sherlock [38] proved that proposal distributions following a geometric “shell” condition from Assumption 2.2 (6) should yield an optimal acceptance rate of approximately 0.234 when sampling from a suitable target. As both the Laplace and Uniform proposals satisfy this condition, the 0.234 rule should hold for high dimensions, and our experiments serve to verify this theoretical prediction in lower dimensions.
@@
$Q_{L}(\mathbf{y}\mid\mathbf{x})\propto\prod_{i=1}^{d}\exp\left(-\frac{|x_{i}|}{\sigma}\right),\quad Q_{U}(\mathbf{y}\mid\mathbf{x})=\prod_{i=1}^{d}\frac{\mathbb{I}\left[|y_{i}-x_{i}|\leq\tfrac{b}{2}\right]}{b}$
The first proposal distribution we consider is the multivariate Laplace distribution, which may also be referred to as the double exponential distribution. We use location parameter $\mu=0$ and covariance parameter $\Sigma=\sigma^{2}I_{d}$, with $\sigma$ to be chosen. We also consider an example with the uniform proposal distribution within the closed interval $\left[x_{i}-\frac{b}{2},x_{i}+\frac{b}{2}\right]$ for each component $i$, where $b>0$ is the interval length.
@@
These empirical findings provide strong validation for the generalized theory of optimal scaling. While the foundational result of Roberts et al. [29] was derived for Gaussian proposals, the more general criteria from Sherlock [38], summarized in Assumption 2.2, extend this theory. The key requirement on the proposal is the shell condition (6) that requires the norm of the proposal vector to concentrate around a constant in high dimensions. Both the multivariate Laplace and Uniform distributions used here are spherically symmetric and satisfy this condition. The theory therefore predicts that the 0.234 acceptance rate should indeed be asymptotically optimal when using these proposals on a suitable target. Our results confirm this prediction, and characterize the finite-dimensional behaviour, demonstrating that convergence to the 0.234 optimum occurs from above for these proposals.

### v4

@@ Abstract
For random-walk Metropolis (RWM) and parallel tempering (PT) algorithms, an asymptotic acceptance rate of around 0.234 is known to be optimal in certain high-dimensional limits. However, its practical relevance is uncertain due to restrictive derivation conditions. We synthesise previous theoretical advances in extending the 0.234 acceptance rate to more general settings, and demonstrate its applicability with a comprehensive empirical simulation study on examples examining how acceptance rates affect Expected Squared Jumping Distance (ESJD). Our experiments show the optimality of the 0.234 acceptance rate for RWM is surprisingly robust even in lower dimensions across various non-spherically symmetric proposal distributions, multimodal target distributions that may not have an i.i.d. product density, and curved Rosenbrock target distributions with nonlinear correlation structure. Parallel tempering experiments also show that the idealized 0.234 spacing of inverse temperatures may be approximately optimal for low dimensions and non i.i.d. product target densities, and that constructing an inverse temperature ladder with spacings given by a swap acceptance of 0.234 is a viable strategy.
@@ 1 Introduction
[30] proved an important result for independently and identically distributed (i.i.d.) product target densities: in the state space high-dimensional limit $d\to\infty$, the asymptotic acceptance rate of a Metropolis algorithm is central to its efficiency. Under their assumptions, an asymptotic acceptance rate of approximately 0.234 maximises the efficiency of a random-walk Metropolis (RWM) algorithm with a Gaussian proposal distribution starting at stationarity. However, their assumptions on the target and proposals are very restrictive. Many practical scenarios often involve finite and relatively small dimensions. Additionally, many probability distributions do not have densities of the i.i.d. product form, and for those that do, sampling a one-dimensional target is much simpler than using RWM. A burgeoning active research area is dedicated to showing that this figure still applies beyond these assumptions, and much of this focuses on weakening the i.i.d. product assumption.
Many papers have shown that the optimal average acceptance rate of 0.234 for RWM holds in greater generality than the assumptions used in Roberts et al. [30]. For more general theoretical results, Sherlock [39] replaces the restrictive i.i.d. product target and Gaussian proposal setting of Roberts et al. [30] with a set of “shell”, relative variability, and eccentricity conditions on the target, eccentricity condition on the jump distance matrix, and shell condition on the proposal. This framework does not require a product form and is based on the geometric properties of the (log) target distribution in high dimensions. Yang et al. [43] makes the 0.234 result applicable to targets with sparse dependency structures, such as those arising from Bayesian graphical models, using a set of sufficient conditions based on local dependencies and bounded asymptotic behaviour of derivatives. We expand on these results in Section 2. The 0.234 acceptance rate in optimal scaling has also been shown to apply to the spacing and swapping between inverse temperatures in the parallel tempering MCMC method [1, 37]; we describe this in greater detail in Section 4.
That being said, there are some cases in which 0.234 is not optimal, such as in the case of discontinuous targets [26] where the optimal acceptance rate is approximately 0.1353. Additionally, these results of optimality are still theoretical ideals, with required assumptions that could be further relaxed, and in the limit of $d\to\infty$, where the Markov chain converges to a diffusion process.
@@
These results are of great interest to practitioners who would like to understand how to tune their MCMC algorithms to maximise their efficiency. Previous simulation studies have been carried out on more general target distributions like products of independent, non-identical components [8, 34], and for the Markov modulated Poisson process [38], but there is still a lack of thorough experimentation and guidance for how applicable the 0.234 rule is on realistic target distributions that may have lower or higher dimensionality and/or multimodality. Our research examines how necessary these assumptions are for the theoretical result to remain relevant: where can we relax some assumptions in lower dimensions and still have the optimal acceptance rate of approximately 0.234, and where can we not do this? And in the end, how applicable is the 0.234 figure for the practitioner, who may want to use MCMC for complicated target distributions beyond the necessary assumptions in finite dimensions? We thoroughly dissect various aspects of the Metropolis algorithm and experiment with them, and in doing so, we show empirically where the theoretical ideal value can still align with more realistic scenarios.
@@ 2.1 Metropolis Algorithm
A Metropolis algorithm [22] constructs a Markov chain $\mathbf{x}^{(0)},\mathbf{x}^{(1)},\dots,\mathbf{x}^{(t)}$. If the chain is constructed to be irreducible and aperiodic, it is guaranteed to have the target distribution $\pi$ as its unique stationary distribution. This means that as the number of steps $t$ becomes large, the distribution of $\mathbf{x}^{(t)}$ converges to $\pi$. Therefore, after an initial burn-in period, the subsequent states of this single, long chain can be used as a sample from the target distribution $\pi$. When taking each step in the Markov chain, we generate a new state $\mathbf{y}$ from a proposal distribution $Q(\cdot|\mathbf{x}^{(t)})$ that is conditional on the current state $\mathbf{x}^{(t)}$ and is symmetric, meaning $Q(\mathbf{y}|\mathbf{x}^{(t)})=Q(\mathbf{x}^{(t)}|\mathbf{y})$. We accept this new state with probability $\alpha(\mathbf{x},\mathbf{y})$ given by $\alpha=\min\left\{1,\dfrac{\pi(\mathbf{y})}{\pi(\mathbf{x})}\right\}$, where, by a common abuse of notation, we let $\pi(\cdot)$ also denote the probability density function of the target distribution $\pi$. If the proposed value is accepted, we set the next state $\mathbf{x}^{(t+1)}=\mathbf{y}$. Otherwise, we set $\mathbf{x}^{(t+1)}=\mathbf{x}^{(t)}$.
@@ 2.2 Optimal Scaling Framework
Optimal Scaling [30, 34] is one of the most successful frameworks for performing asymptotic analysis of high-dimensional MCMC methods, and provides mathematically-grounded guidance on how to best optimise MCMC performance by tuning the “scaling” parameter(s) of the proposal distribution. Typically, a multivariate Gaussian distribution centred at the current state is used as the proposal distribution. The scaling of the Gaussian proposal is then given by a scaling factor $\sigma$, and this is often used as a variance $\sigma^{2}$ that is applied to the $d\times d$ identity matrix $I_{d}$. So, $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{d})$. If the proposal’s variance is too small, many proposed steps are accepted by the algorithm, but each step does not explore the state space much. Vice versa, if the proposal’s variance is too large, too few proposed steps are accepted, leading to slow exploration of the state space since the algorithm stays at a state for too long. Therefore, the proposal’s variance is crucial to the performance of the algorithm; the optimal scaling should balance the exploration of new areas in the state space with exploiting high-density areas of the target distribution.
@@
where $n$ is the total number of iterations of the algorithm. Maximising the ESJD aligns with minimizing the first-order auto-correlation of the Markov chain and subsequently maximises efficiency if each higher order auto-correlation is a monotonically increasing function of the first-order auto-correlation [27, 43]. It is worth mentioning that ESJD primarily measures local convergence performance. Because heavy-tailed targets disrupt the monotonic relationship between first-order and higher-order autocorrelations, achieving global convergence for heavy-tailed targets requires maximising an adapted version of ESJD [18]. However, we do not consider heavy-tailed targets in this work, so the working definition of ESJD is sufficient.
There are other notions of efficiency of a Markov chain, but in the high-dimensional limit $d\to\infty$, if the moment conditions in Assumption 2.1 hold, the chain converges to a diffusion process, and all efficiency measures are effectively equivalent [30, 34]. A key result of these referenced papers is that, given a Metropolis algorithm with a Gaussian proposal distribution $\mathcal{N}(\mathbf{x}^{(t)},\frac{\ell^{2}}{d}I_{d})$ where $\ell>0$ is a fixed scaling constant and $I_{d}$ is the identity matrix, maximising the diffusion’s speed measure $h(\ell)$, which is a function of the scaling constant $\ell$, yields the most efficient asymptotic diffusion. Furthermore, the speed measure has a clear relation to a much simpler quantity to tune for: the asymptotic acceptance rate of the proposed new states (moves) of the algorithm, defined as
@@
Under the conditions in Assumption 2.1, Roberts et al. [30] showed that for an RWM algorithm with a Gaussian proposal distribution starting in stationarity, when the whole process is sped-up by a factor of $d$, the first component of the process converges to a limiting Langevin diffusion. Maximizing the speed of this limiting process, in turn, corresponds to tuning the algorithm to an asymptotic acceptance rate of approximately 0.234.
@@
(Sherlock [39]) Let $\mathbf{x}$ be a random variable from the target density $\pi$, $M(\mathbf{x})=||\nabla\log\pi(\mathbf{x})||$ denote the norm of the gradient of the log target, $H(\mathbf{x})=-\nabla^{2}\log\pi(\mathbf{x})$ denote the negative Hessian matrix of the log target, and $\Delta(\mathbf{x},\mathbf{u})=H(\mathbf{x}+\mathbf{u})-H(\mathbf{x})$ denote the matrix of perturbations of the negative Hessian. The primary conditions are:
@@
Put together, these geometric regularities recover the same result of the optimality of the 0.234 acceptance rate as Roberts et al. [30] but with weaker assumptions on the target and proposal.
@@ 3 Random-Walk Metropolis Simulations
In this section, we explore the applicability of the 0.234 acceptance rate figure for RWM algorithms using several practical RWM examples in lower dimensions. We begin with simple examples where the target densities meet all the key conditions in Roberts et al. [30] (provided as Assumption 2.1) required for the 0.234 acceptance rate to be optimal. Next, we investigate examples that do not use Gaussian proposal distributions, and then we explore other target densities that may lack these assumptions, such as smoothness or the i.i.d. product form. We estimate the Expected Squared Jumping Distance (ESJD) of the algorithm as a metric of the algorithm’s efficiency as mentioned in Section 2.2, and aim to see under what conditions an acceptance rate of approximately 0.234 maximises the ESJD for various target and proposal distributions in lower dimensions.
@@
The first example sets $f_{1}$ with shape parameter 3 and scale parameter 2, and the second example sets $f_{2}$ with shape parameters 3 and 2. We use the multivariate Gaussian proposal distribution defined in (8) in both examples. We run this experiment in dimensions $d\in\{2,5,10,30,50,100\}$.
@@
Since most ESJD versus acceptance rate plots look fairly similar and have a similar progression as dimension changes, the full set of curves for ESJD versus acceptance rate plots for other subsections are in Appendix A, including the results for the i.i.d. Gamma in Figure 9. As an aside, in all dimensions, there is a relatively broad plateau where the acceptance rate is approximately optimal; thus, the coarse 40-point grid pinpoints the optimum to approximately ±0.01$\pm$ 0.01± 0.01.
@@ 3.2 Non-Gaussian Proposal Densities
Next, we investigate the efficacy of different proposal distributions for the Metropolis algorithm. One of the conditions assumed by the 0.234 theorem of Roberts et al. [30] is the use of a multivariate Gaussian proposal distribution $Q(\mathbf{y}|\mathbf{x})=\mathcal{N}(\mathbf{y}|\mathbf{x},\sigma^{2}I_{d})$. The additional results in [34] depend on the assumptions of light-tailed proposals. In contrast, [17] extend this analysis by examining the behaviour of the algorithms with heavy-tailed proposals. They find that, unlike the diffusion behaviour seen with light-tailed proposals, heavy-tailed proposals lead to a different dynamic characterized by abrupt movements followed by periods of inactivity. Furthermore, [25] investigated optimal scaling results for heavy-tailed proposal distributions such as the Cauchy distribution, and spherical proposal distributions of fixed radius. In their RWM simulation study with three different continuous i.i.d. product target densities, both the Gaussian and Cauchy proposals had an estimated asymptotically optimal acceptance rate of approximately 0.234 in $d=100$, and a spherical proposal with a fixed radius of 2.38 achieved an optimal acceptance rate of about 0.234 in $d=2$ and $d\geq 10$. Finally, Sherlock [39] proved that spherically symmetric proposal distributions following a geometric “shell” condition from Assumption 2.2 (6) should yield an optimal acceptance rate of approximately 0.234 when sampling from a suitable target.
@@
$Q_{L}(\mathbf{y}\mid\mathbf{x})\propto\prod_{i=1}^{d}\exp\left(-\frac{|y_{i}-x_{i}|}{\sigma}\right),\quad Q_{U}(\mathbf{y}\mid\mathbf{x})=\prod_{i=1}^{d}\frac{\mathbb{I}\left[|y_{i}-x_{i}|\leq\tfrac{b}{2}\right]}{b}$
The first proposal distribution we consider is the multivariate Laplace distribution, which may also be referred to as the double exponential distribution. We set the location parameter $\mu=\mathbf{x}$ and scale parameter $\sigma$ to be chosen to influence the acceptance rate of the algorithm. We also consider an example with the uniform proposal distribution within the closed interval $\left[x_{i}-\frac{b}{2},x_{i}+\frac{b}{2}\right]$ for each component $i$, where $b>0$ is the interval length.
@@
These empirical findings provide strong validation for the generalized theory of optimal scaling. While the foundational result of Roberts et al. [30] was derived for Gaussian proposals, the more general criteria from Sherlock [39], summarized in Assumption 2.2, extend this theory. The key requirement on the proposal is the shell condition (6) that requires the norm of the proposal vector to concentrate around a constant in high dimensions. Interestingly, although both the multivariate Laplace and Uniform distributions used here are not spherically symmetric as required by (6), they still satisfy the spirit of the shell condition mentioned in the previous sentence. If these more relaxed conditions still satisfy the necessary requirements for the theory to work, the theory predicts that the 0.234 acceptance rate should indeed be asymptotically optimal when using these proposals on a suitable target. Our results confirm this prediction, and characterize the finite-dimensional behaviour, demonstrating that convergence to the 0.234 optimum occurs from above for these proposals. The lack of spherical symmetry in our proposals also suggests that weaker conditions on the proposal than spherical symmetry may be sufficient for the 0.234 rate to be optimal.